OpenStudy (anonymous):

What is the simplified form of 24 y to the fifth power over 15 x to the eighth power divided by 8 y squared over 4 x to the fourth power ? 4 y cubed over 5 x to the fourth power 4 y to the fourth power over 5 x cubed 5 x to the fourth power over 4 y cubed 5 x cubed over 4 y to the fourth powerus 3y

OpenStudy (anonymous):

\[\frac{ 24y ^{5} }{ 15x ^{8} } \div \frac{ 8y ^{2} }{ 4x ^{4} }\]

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

We can flip the 2nd fraction and multiply..

OpenStudy (igreen):

\(\sf \dfrac{24y^5}{14x^8} \div \dfrac{8y^2}{4x^4} \rightarrow \dfrac{24y^5}{14x^8} \times \dfrac{4x^4}{8y^2}\)

OpenStudy (igreen):

Then multiply the numerators and denominators together. \(\sf \dfrac{24y^5}{14x^8} \times \dfrac{4x^4}{8y^2} \rightarrow \dfrac{24y^5 \times 4x^4}{14x^8 \times 8y^2}\)

OpenStudy (anonymous):

okay then what? do u have to cross multiply or horizontally multiply

OpenStudy (igreen):

I just showed you, we do it horizontally..we only cross multiply when we have an equal sign between the fractions.

OpenStudy (igreen):

Can you simplify \(\sf \dfrac{24y^5 \times 4x^4}{14x^8 \times 8y^2}\) ?

OpenStudy (anonymous):

yeah but non of those are my nswer choies can you simplify it even further if possible?

OpenStudy (anonymous):

yeah i think so

OpenStudy (anonymous):

but how would you simplify it further because i know you can reduce it

OpenStudy (igreen):

Hold on

OpenStudy (anonymous):

k

OpenStudy (igreen):

Okay, first we divide the numbers on the second fraction, \(\sf \dfrac{8y^2}{4x^4}\).

OpenStudy (igreen):

8 / 4 = ?

OpenStudy (anonymous):

2

OpenStudy (igreen):

Yes, so we have \(\sf\dfrac{2y^2}{x^4}\).

OpenStudy (igreen):

Now simplify \(\sf\dfrac{24}{15}\) from \(\sf\dfrac{24y^5}{15x^8}\).

OpenStudy (anonymous):

8y^5 / 3x^8

OpenStudy (igreen):

Check again

OpenStudy (igreen):

\(\sf 8\) is correct, but \(\sf 3\) is not..

OpenStudy (anonymous):

would it be 5

OpenStudy (igreen):

Yes.

OpenStudy (anonymous):

ok

OpenStudy (igreen):

So we have: \(\sf \dfrac{8y^5}{5x^8} \div \dfrac{2y^2}{x^4}\)..now we cross multiply(I was wrong earlier..lol)

OpenStudy (anonymous):

okay np give me a sec o do that plz

OpenStudy (anonymous):

would it be 8y^5 x^4 / 5x^8 2y^2

OpenStudy (anonymous):

but if we simplify it would the final answer be \[\frac{ 4y^3 }{ 5x^4 }\]

OpenStudy (igreen):

Yep, you got it.

OpenStudy (anonymous):

thnx